# A machine-learning classifier for the postmerger remnant of binary neutron stars

### v1

@@ Abstract
Knowing the kind of remnant produced after the merger of a binary neutron star system, e.g., if a black hole forms or not, would not only shed light on the equation of state describing the extremely dense matter inside neutron stars, but also help understand the physical processes involved in the postmerger phase. Moreover, in the event of a gravitational-wave detection, predicting the presence of a neutron star remnant is crucial in order to advise potential electromagnetic follow-up campaigns. In this work, we use Gradient Boosted Decision Trees and publicly available data from numerical-relativity simulations to construct a classifier that predicts the outcome of binary neutron star mergers, based on the binary’s parameters inferred from gravitational-wave inspiral signals: total mass, mass-weighted tidal deformability, mass ratio, and effective inspiral spin. Employing parameters that can be estimated from the inspiral part of the signal only allows us to predict the remnant independently on the detection of a postmerger gravitational-wave signal. We build three different classifiers to distinguish between various potential scenarios, we estimate their accuracy and the confidence of their predictions. Finally, we apply the developed classifiers to real events data, finding that GW170817 most likely lead to the formation of a hypermassive neutron star, while GW190425 to a prompt collapse to a black hole.
@@ I Introduction
Binary neutron star (BNS) mergers allow us to study matter at extremely high densities (see, e.g., Refs. [1, 2, 3] for a review). One of the main tools to investigate these phenomena is the gravitational-wave (GW) signal emitted during the entire coalescence, since the GWs morphology strongly depends on the equation of state (EoS) of the matter forming the neutron stars (NSs) in the system [4]. Advanced LIGO [5] and Advanced Virgo [6] already observed two such events, GW170817 [7] and GW190425 [8], and more detections are foreseen in the upcoming years, both with current and especially with future-generation detectors, such as the Einstein Telescope [9, 10, 11, 12] and the Cosmic Explorer [13, 14, 15].
@@
  • The lifetime of the remnant plays a crucial role in the potential electromagnetic counterparts; cf. Ref. [50] and references therein.
@@ II Methods
Figure 1: Schematic representation of the possible outcomes of a BNS merger depending on the system’s total mass. We show the different kinds of remnants with grayscale bars that gradually progress towards black as the system evolves towards the final collapse to a BH. The red vertical dashed lines represent the thresholds on the remnant lifetime employed to distinguish the different classes for each classifier. The colored boxes (orange for Classifier A, teal for Classifier B, and pink for Classifier C) show how the various remnants are grouped in the different classes employed for each classifier.
In this section, we describe how we build the dataset, with a focus on the procedure employed to split the data points in training and validation set. We briefly illustrate how the GBDT algorithm works and explain how we evaluate the performance of our models.
@@ II.1 Dataset
To build our dataset, we employ numerical-relativity (NR) simulations data from the CoRe [65, 66] and SACRA [67] databases, together with simulations from Refs. [68, 69]. The simulations considered span a wide range of values for the binary system’s parameters and employ different EoSs, including also phase transitions. When the remnant collapse time is not directly provided, we compute it from the GW data, as the difference between the merger time and the time when the GW signal goes to zero (see bottom panel of Fig. 2). If the GW signal does not go to zero, we register the remnant lifetime as “greater” than the simulation time after the merger, computed as the difference between the merger time and the total simulation time, after removing the last 2 ms to take into account the effects of data tapering. Regarding the data in Refs. [68, 69], for the configurations that do not collapse within simulation time, we consider the remnant lifetime to be $>25\,{\rm ms}$, where $25\,{\rm ms}$ is a conservative estimate of the duration of these simulations after the merger. If the system total mass is lower than $M_{\rm\textsc{tov}}$ for the specific EoS employed in the simulation, we classify the remnant as “stable”. For each system, we consider data from the simulation with the highest resolution and the largest extraction radius available [^1] More details about the selected simulations can be found at [64].. We discard eccentric configurations, and, among the simulations in Ref. [69], we do not include the ones for misaligned-spin configurations. In the end, our dataset includes 398 systems.
@@
    Classifier C discriminates between all the four different scenarios listed above; see Sec. III.3.
@@ II.2 Data splitting
Moreover, the performance of the classifier is strongly influenced by the random seed utilized in the training-validation splitting. This is due to the fact that some data points are more difficult to classify, for various reasons (for example, when $\tau_{\rm BH}$ is close to the threshold values that determine the class, or when some configurations have the same parameters but different spin values and can therefore result in different outcomes). This effect can be mitigated by ensuring that at least some of the points with “problematic” features are included in the training set, in order to improve their modeling. On the other hand, we also want to include some of them in the validation set, to make sure that the evaluation of the classifier performance is not carried out only on “easy” points and thus biased. To accomplish both these goals, we consider training-validation splittings with 1000 different seeds: for each one of them, we train a GBDT model with default parameters and fixed random state [^2] Optimizing the hyperparameters for each model would decrease the number of misclassified points, but at the expense of a much larger computational cost. Hence, we keep the default parameters, with the resulting misclassifications representing a conservative scenario., and save the misclassified points as “bad” points. We further divide the final batch of “bad” points in three groups based on the number of their misclassification occurrences: less than 10 times, between 10 and 90 times, and more than 90 times, with the cutoff values chosen arbitrary. The final training (validation) set is built by taking $90\%$ ($10\%$) of the points that were never misclassified, together with $90\%$ ($10\%$) of the points in each one of the “bad” points groups. With this procedure to include “bad” points both in the training and validation set, we consider again different splittings created with different random seeds, and, to train and test the classifier, we choose among these one of the training-validation splittings yielding the highest validation accuracy. This approach ensures that the model is trained to recognize also the difficult points, and that the final evaluation for the model performance on the validation set is not biased.
@@ II.3 Gradient Boosted Decision Trees Classifier
Decision trees are a non-parametric supervised learning technique, whose output is based on a series of hierarchical decision rules inferred from the data [74]. In ensemble methods, like random forests or boosting methods [75, 76, 77], multiple decision trees are combined, in order to improve the robustness of the predictions. In Gradient Boosted Decision Trees, the prediction for an input $\mathbf{x}_{i}$ is given by
@@
where the weak learners $h_{m}$ are single decision trees, and the ensemble model $F_{\textsc{m}}$ is also called the strong model [^3] Decision trees predictions are continuous values, hence, in classification problems, the final prediction, i.e., the discrete value corresponding to a class, needs to be computed from the continuous $F_{\textsc{m}}(\mathbf{x}_{i})$ with a function depending on the loss function employed. In the case of a log-loss error, a sigmoid and softmax functions are used for binary and multiclass classification problems, respectively [77].. The gradient boosting model is built through an iterative approach [77], with the goal to minimize the expectation value of the loss function $L$, which in our case is a log-loss, on the data set $\left\{\mathbf{x}_{i},y_{i}\right\}_{i=1}^{N}$. At each step $t$, the strong model is updated with a new weak learner
@@
where the shrinkage ratio (or learning rate in sklearn) $\epsilon\in[0,1]$ is a parameter introduced to reduce overfitting. The weak learner is chosen from a family of functions $H$ in order to minimize the average of the loss function on the dataset of the strong learner at the previous step
@@
The performance of GBDT models can be improved by tuning their hyperparameters (see, e.g., Ref. [78]). Regarding the shrinkage ratio $\epsilon$, although larger values imply that the model learns faster, smaller values help avoid overfitting; typically, we employ $\epsilon\sim 0.1$. Another useful parameter is the total number $M$ of weak learners comprising the model (dubbed n_estimators in sklearn): to avoid overly complex models, an adequate value of $M$ can be estimated via early stopping [79, 80, 81, 82]. Finally, the complexity and size of each weak learner is controlled by the max_depth parameter, which limits the number of nodes in each tree. In GBDT, the component trees are usually quite shallow, and we use max_depth $<5$.
@@ II.4 Performance of the classifier
where $\hat{y}_{i}$ is the predicted label for the $i$-th sample, $y_{i}$ its true label, $N$ the total number of samples, and $\mathbf{1}$ the indicator function. The accuracy, however, does not take into account the distribution of correct and wrong predictions across the different classes, and the information it provides can be misleading in the case of highly imbalanced datasets.
Let us consider the case of a binary classification, where usually the predictions are labeled as “positive” (or 1, which in our case could be the presence of a NS remnant) and “negative” (or 0, in our case the prompt collapse to BH). The confusion matrix obtained with our model on a given set of data will include four different kinds of entries: true positives (TP), the data points correctly classified as positive, false positives (FP), data whose true label is negative but are misclassified as positive, true negatives (TN), points correctly classified as negative, and false negatives (FN), positives misclassified as negative. In the binary case, the accuracy is
@@ III Results
In this section, we discuss the three classifiers, both in terms of their accuracy and predictions on the two real events GW170817 and GW190425. For Classifier A, we also investigate the influence of the binary parameters in determining whether the system will promptly collapse to a BH or not.
@@ III.1 Classifier A
As a first step, we build a classifier to distinguish between prompt collapse and the formation of a NS remnant. We divide the 398 points in a training and a validation set as explained in Sec. II.2, and then we optimize the hyperparameters of the model using GridSearchCV [85], manual tuning, and early stopping.
@@
The sklearn GBDT algorithm provides also the features importances for the model, computed as the Gini importance, and we find that the parameter that affects the classification the most is $\tilde{\Lambda}$ ($\sim 0.60$), followed by $M_{\rm tot}$ ($\sim 0.20$), $q$ ($\sim 0.14$), and finally $\chi_{\rm eff}$ ($\sim 0.04$). Given the possible inconsistency among values of the features’ importance computed with different methods (e.g., gain, split count, or permutation method) [86], a new measure of the contribution of each feature to a model prediction has been proposed, the SHAP (SHapley Additive exPlanations) values [87]. The SHAP method derives from game theory [88] and belongs to the Explainable Artificial Intelligence research field [89], which investigates possible methods to explain models obtained with artificial intelligence tools that would otherwise be considered “black-boxes”. The SHAP technique assigns locally, i.e., for each data point, an importance to each feature based on its impact on the model’s prediction for that point. For Classifier A, we compute SHAP values for the points in the training set with the shap Python package [87, 86, 90] [^4] For the moment, the computation of SHAP values for GBDT models is supported only for binary classifiers.. The SHAP values are related to the log odds predicted by the model: more positive (negative) SHAP values “push” the model prediction to larger positive (negative) values, which, for Classifier A, translate to a larger probability for the prompt collapse to BH (production of a NS remnant) scenario. Therefore, to evaluate the importance of a feature, one needs to look at the absolute value of its SHAP measure. Computing the mean absolute SHAP values for each feature, we find again that the parameter with the largest impact on the model prediction is $\tilde{\Lambda}$, followed by $M_{\rm tot}$, $q$, and $\chi_{\rm eff}$. However, if we look at the maximum of the SHAP absolute values, the second most important feature is the mass ratio $q$. This means that, on average, $M_{\rm tot}$ has a stronger impact than $q$, but, for the single data points, $q$ can actually affect the prediction more than $M_{\rm tot}$. Figure 4 shows the summary plot for the computed SHAP values. In addition to the order of the features based on their importance, which we already discussed, we see that, as expected, higher values for $M_{\rm tot}$ are more likely to push the prediction towards prompt collapse and the same holds for large $q$ values. The effective spin $\chi_{\rm eff}$ is clearly the feature with the more modest impact, and indeed most of the points are clustered around very small absolute values of SHAP. Although, on average, the SHAP values are small, for a few points, the impact of $\chi_{\rm eff}$ is non-negligible. We notice that negative spin values tend to influence the prediction towards a prompt collapse, while positive ones towards the production of an NS remnant; we will discuss the spin effect in more detail later.
@@
We also note that only a limited number of NR simulations for spinning systems are available, thus the very weak contribution of spin to the overall classification might not necessarily be interpreted as a weak influence of this parameter on the remnant fate, but rather as the need for more spinning simulations to better understand the role of such parameter; cf. works in Ref. [38, 69].
@@
Table 2: Probabilities for the different merger outcomes for the three classifiers. We remind here that Classifier A discriminates between prompt collapse to BH ($p_{\rm\textsc{pcbh}}$) and formation of a NS remnant ($p_{\rm\textsc{rns}}$); Classifier B distinguishes between prompt collapse ($p_{\rm\textsc{pcbh}}$), formation of a HMNS that collapses during the simulation ($p_{\rm\textsc{hmns}}$), and formation of a NS remnant (which could be a HMNS, SMNS or stable NS) that does not collapse within simulation time ($p_{\rm\textsc{nc}}$); finally, Classifier C includes four classes: prompt collapse to BH ($p_{\rm\textsc{pcbh}}$), production of a HMNS that collapses during the simulation, with $2\,{\rm ms}<\tau_{\rm BH}<5\,{\rm ms}$ ($p_{\rm\textsc{short}}$) or $\tau_{\rm BH}>5\,{\rm ms}$ ($p_{\rm\textsc{long}}$), and formation of a remnant that does not collapse within simulation time ($p_{\rm\textsc{nc}}$).

### v2

@@ Abstract
Knowing the kind of remnant produced after the merger of a binary neutron star system, e.g., if a black hole forms or not, would not only shed light on the equation of state describing the extremely dense matter inside neutron stars, but also help understand the physical processes involved in the postmerger phase. Moreover, in the event of a gravitational-wave detection, predicting the presence of a neutron star remnant is crucial in order to advise potential electromagnetic follow-up campaigns. In this work, we use Gradient Boosted Decision Trees and publicly available data from numerical-relativity simulations to construct a classifier that predicts the outcome of binary neutron star mergers, based on the binary’s parameters inferred from gravitational-wave inspiral signals: total mass, mass-weighted tidal deformability, mass ratio, and effective inspiral spin. Employing parameters that can be estimated from the inspiral part of the signal only allows us to predict the remnant independently on the detection of a postmerger gravitational-wave signal. We build three different classifiers to distinguish between various potential scenarios, we estimate their accuracy and the confidence of their predictions. Finally, we apply the developed classifiers to real events data, finding that GW170817 most likely led to the formation of a hypermassive neutron star, while GW190425 to a prompt collapse to a black hole.
@@ I Introduction
Binary neutron star (BNS) mergers allow us to study matter at extremely high densities (see, e.g., Refs. [1, 2, 3] for a review). One of the main tools to investigate these phenomena is the gravitational-wave (GW) signal emitted during the entire coalescence, since the GWs morphology strongly depends on the equation of state (EoS) of the matter forming the neutron stars (NSs) in the system (see, e.g., Refs. [4, 5, 1, 6, 7, 8] and references therein). Advanced LIGO [9] and Advanced Virgo [10] already observed two such events, GW170817 [11] and GW190425 [12], and more detections are foreseen in the upcoming years, both with current and especially with future-generation detectors, such as the Einstein Telescope [13, 14, 15, 16] and the Cosmic Explorer [17, 18, 19].
@@
  • The lifetime of the remnant plays a crucial role in the potential electromagnetic (EM) counterparts; cf. Ref. [54] and references therein.
@@ II Methods
Figure 1: Schematic representation of the possible outcomes of a BNS merger depending on the system’s total mass. We show the different kinds of remnants with grayscale bars that gradually progress toward black as the system evolves toward the final collapse to a BH. The red vertical dashed lines represent the thresholds on the remnant lifetime employed to distinguish the different classes for each classifier. The colored boxes (orange for Classifier A, teal for Classifier B, and pink for Classifier C) show how the various remnants are grouped in the different classes employed for each classifier.
In this section, we describe how we build the dataset, with a focus on the procedure employed to split the data points into training and validation sets. We briefly illustrate how the GBDT algorithm works and explain how we evaluate the performance of our models.
@@ II.1 Dataset
To build our dataset, we employ numerical-relativity (NR) simulations data from the CoRe [69, 70] and SACRA [71] databases, together with simulations from Refs. [72, 73]. The simulations considered span a wide range of values for the binary system’s parameters and employ different EoSs, also including phase transitions. When the remnant collapse time is not directly provided, we compute it from the GW data, as the difference between the merger time and the time when the GW signal goes to zero (see the bottom panel of Fig. 2). If the GW signal does not go to zero, we register the remnant lifetime as “greater” than the simulation time after the merger, computed as the difference between the merger time and the total simulation time, after removing the last 2 ms to take into account the effects of data tapering. Regarding the data in Refs. [72, 73], for the configurations that do not collapse within the simulation time, we consider the remnant lifetime to be $>25\,{\rm ms}$, where $25\,{\rm ms}$ is a conservative estimate of the duration of these simulations after the merger. If the system total mass is lower than $M_{\rm\textsc{tov}}$ for the specific EoS employed in the simulation, we classify the remnant as “stable”. For each system, we consider data from the simulation with the highest resolution and the largest extraction radius available [^1] In the case of a few configurations that were included both in the CoRe database and in Refs. [72, 73], we kept the collapse time value provided in Refs [72, 73], which is computed directly from the NR data describing the system evolution. More details about the selected simulations can be found at [68].. We discard eccentric configurations, and, among the simulations in Ref. [73], we do not include the ones for misaligned-spin configurations. In the end, our dataset includes 398 systems.
@@
    Classifier C discriminates between all four different scenarios listed above; see Sec. III.3.
@@ II.2 Data splitting
Moreover, the performance of the classifier is strongly influenced by the random seed utilized in the training-validation splitting. This is due to the fact that some data points are more difficult to classify, for various reasons (for example, when $\tau_{\rm BH}$ is close to the threshold values that determine the class, or when some configurations have the same parameters but different spin values and can, therefore, result in different outcomes). This effect can be mitigated by ensuring that at least some of the points with “problematic” features are included in the training set, in order to improve their modeling. On the other hand, we also want to include some of them in the validation set, to make sure that the evaluation of the classifier performance is not carried out only on “easy” points and thus biased. To accomplish both these goals, we create 1000 distinct training-validation splittings, using a different seed for each one. For each splitting, we train and validate a GBDT model with a fixed random state and with the default parameters provided for the GradientBoostingClassifier class in sklearn [^2] Optimizing the hyperparameters for each model would decrease the number of misclassified points, but at the expense of a much larger computational cost. Hence, we keep the default parameters, with the resulting misclassifications representing a conservative scenario. On the other hand, the random state, which controls the randomness present in various steps of the algorithm, is not fixed by default in the sklearn model. We specify it not only to ensure reproducibility, but also to avoid possible additional sources of uncertainty, since at this stage we want to focus on the variations induced by the different training-validation splittings. We arbitrarily set random_state$=42$; for completeness, we checked the effect of different values for the GBDT model random state and found that results are not significantly affected., and save the misclassified points as “bad” points. We further divide the final batch of “bad” points in three groups based on the number of their misclassification occurrences: less than 10 times, between 10 and 90 times, and more than 90 times, with the cutoff values chosen arbitrary. The final training (validation) set is built by taking $90\%$ ($10\%$) of the points that were never misclassified, together with $90\%$ ($10\%$) of the points in each one of the “bad” points groups. With this procedure to include “bad” points in both the training and validation sets, we consider again different splittings created with different random seeds, and, to train and test the classifier, we choose among these one of the training-validation splittings yielding the highest validation accuracy. This approach ensures that the model is trained to recognize also the difficult points, and that the final evaluation for the model performance on the validation set is not biased.
Table 1 reports the ranges for parameters in both the training and complete datasets for all three classifiers. We highlight that the distribution of points in our dataset is not uniform in these ranges, since it depends on the available NR simulations.
Classifier A Classifier B Classifier C
Training Total Training Total Training Total
$M_{\rm tot}$ [2.4, 3.4] $M_{\odot}$ [2.4,3.4] $M_{\odot}$ [2.4,3.4] $M_{\odot}$ [2.4, 3.4] $M_{\odot}$ [2.4, 3.4] $M_{\odot}$ [2.4,3.4] $M_{\odot}$
$q$ [1.0, 2.1] [1.0, 2.1] [1.0, 2.0] [1.0,2.0] [1.0, 2.0] [1.0,2.0]
$\tilde{\Lambda}$ [89, 3520] [89, 6255] [89, 3520] [89, 6255] [89, 6255] [89, 6255]
$\chi_{\rm eff}$ [-0.267, 0.409] [-0.267, 0.409] [-0.267, 0.272] [-0.267, 0.272] [-0.267, 0.267] [-0.267, 0.272]
Table 1: Ranges of parameters covered by the training and total dataset for the different classifiers. As will be discussed in Sec. III.2, the dataset employed for Classifier B and Classifier C is different from the dataset for Classifier A, since it does not include simulations that do not collapse within simulation time, for which the simulation time after the merger is $t_{\rm sim}<25$ ms; consequently, the ranges in the “Total” columns are slightly different.
@@ II.3 Gradient Boosted Decision Trees Classifier
Decision trees are a non-parametric supervised learning technique, whose output is based on a series of hierarchical decision rules inferred from the data [78]. In ensemble methods, such as random forests or boosting methods [79, 80, 81], multiple decision trees are combined, in order to improve the robustness of the predictions. In Gradient Boosted Decision Trees, the prediction for an input $\mathbf{x}_{i}$ is given by
@@
where the weak learners $h_{m}$ are single decision trees, $M$ represents the total number of weak learners, and the ensemble model $F_{\textsc{m}}$ is also called the strong model [^3] Decision trees predictions are continuous values, hence, in classification problems, the final prediction, i.e., the discrete value corresponding to a class, needs to be computed from the continuous $F_{\textsc{m}}(\mathbf{x}_{i})$ with a function depending on the loss function employed. In the case of a log-loss error, a sigmoid and softmax functions are used for binary and multiclass classification problems, respectively [81].. The gradient boosting model is built through an iterative approach [81], with the goal to minimize the expectation value of the loss function $L$, which in our case is a log-loss, on the dataset $\left\{\mathbf{x}_{i},y_{i}\right\}_{i=1}^{N}$. At each step $t$, the strong model is updated with a new weak learner
@@
where the shrinkage ratio (or learning rate in sklearn) $\epsilon\in[0,1]$ is a parameter introduced to reduce overfitting. The weak learner is chosen from a family $H$ of decision trees in order to minimize the average of the loss function on the dataset of the strong learner at the previous step
@@
The performance of GBDT models can be improved by tuning their hyperparameters (see, e.g., Ref. [82]). Regarding the shrinkage ratio $\epsilon$, although larger values imply that the model learns faster, smaller values help avoid overfitting; typically, we employ $\epsilon\sim 0.1$. Another useful parameter is the total number $M$ of weak learners comprising the model (dubbed n_estimators in sklearn): to avoid overly complex models, an adequate value of $M$ can be estimated via early stopping [83, 84, 85]; for our classifiers, typical values of $M$ are around 100. Finally, the complexity and size of each weak learner is controlled by the max_depth parameter, which limits the number of nodes in each tree. In GBDT, the component trees are usually quite shallow, and we use max_depth $<5$.
@@ II.4 Performance of the classifier
where $\hat{y}_{i}$ is the predicted label for the $i$-th sample, $y_{i}$ is its true label, $N$ is the total number of samples, and $\mathbf{1}$ is the indicator function. The accuracy, however, does not take into account the distribution of correct and wrong predictions across the different classes, and the information it provides can be misleading in the case of highly imbalanced datasets.
Let us consider the case of a binary classification, where usually the predictions are labeled as “positive” (or 1, which in our case could be the presence of a NS remnant) and “negative” (or 0, in our case the prompt collapse to BH). The confusion matrix obtained with our model on a given set of data will include four different kinds of entries: true positives (TP), the data points correctly classified as positive, false positives (FP), data points whose true label is negative but are misclassified as positive, true negatives (TN), points correctly classified as negative, and false negatives (FN), positives misclassified as negative. In the binary case, the accuracy is
@@ III Results
In this section, we discuss the three classifiers, in terms of both their accuracy and predictions on the two real events GW170817 and GW190425. For Classifier A, we also investigate the influence of the binary parameters in determining whether the system will promptly collapse to a BH or not.
@@ III.1 Classifier A
As a first step, we build a classifier to distinguish between prompt collapse and the formation of a NS remnant. We divide the 398 points into a training and a validation set as explained in Sec. II.2, and then we optimize the hyperparameters of the model using GridSearchCV [88], manual tuning, and early stopping.
@@
The sklearn GBDT algorithm provides also the features importances for the model, computed as the Gini importance, and we find that the parameter that affects the classification the most is $\tilde{\Lambda}$ ($\sim 0.60$), followed by $M_{\rm tot}$ ($\sim 0.20$), $q$ ($\sim 0.14$), and finally $\chi_{\rm eff}$ ($\sim 0.04$). Given the possible inconsistency among values of the features’ importance computed with different methods (e.g., gain, split count, or permutation method) [89], a new measure of the contribution of each feature to a model prediction has been proposed, the SHAP (SHapley Additive exPlanations) values [90]. The SHAP method derives from game theory [91] and belongs to the Explainable Artificial Intelligence research field [92], which investigates possible methods to explain models obtained with artificial intelligence tools that would otherwise be considered “black-boxes”. The SHAP technique assigns locally, i.e., for each data point, an importance to each feature based on its impact on the model’s prediction for that point. For Classifier A, we compute SHAP values for the points in the training set with the shap Python package [90, 89, 93]. [^4] For the moment, the computation of SHAP values for GBDT models is supported only for binary classifiers. The SHAP values are related to the log odds predicted by the model: more positive (negative) SHAP values “push” the model prediction to larger positive (negative) values, which, for Classifier A, translate to a larger probability for the prompt collapse to BH (production of a NS remnant) scenario. Therefore, to evaluate the importance of a feature, one needs to look at the absolute value of its SHAP measure. Computing the mean absolute SHAP values for each feature, we find again that the parameter with the largest impact on the model prediction is $\tilde{\Lambda}$, followed by $M_{\rm tot}$, $q$, and $\chi_{\rm eff}$. However, if we look at the maximum of the SHAP absolute values, the second most important feature is the mass ratio $q$. This means that, on average, $M_{\rm tot}$ has a stronger impact than $q$, but, for the single data points, $q$ can actually affect the prediction more than $M_{\rm tot}$. Figure 4 shows the summary plot for the computed SHAP values. In addition to the order of the features based on their importance, which we already discussed, we see that, as expected, higher values for $M_{\rm tot}$ are more likely to push the prediction toward prompt collapse and the same holds for large $q$ values. The effective spin $\chi_{\rm eff}$ is clearly the feature with the more modest impact, and indeed most of the points are clustered around very small absolute values of SHAP. Although, on average, the SHAP values are small, for a few points, the impact of $\chi_{\rm eff}$ is non-negligible. We notice that negative spin values tend to influence the prediction toward a prompt collapse, while positive ones toward the production of an NS remnant; we will discuss the spin effect in more detail later.
@@
We also note that only a limited number of NR simulations for spinning systems are available, thus, the very weak contribution of spin to the overall classification might not necessarily be interpreted as a weak influence of this parameter on the remnant fate, but rather as the need for more spinning simulations to better understand the role of such parameter; cf. works in Ref. [42, 73].
@@
Table 3: Probabilities for the different merger outcomes for the three classifiers. We remind here that Classifier A discriminates between prompt collapse to BH ($p_{\rm\textsc{pcbh}}$) and formation of a NS remnant ($p_{\rm\textsc{rns}}$); Classifier B distinguishes between prompt collapse ($p_{\rm\textsc{pcbh}}$), formation of a HMNS that collapses during the simulation ($p_{\rm\textsc{hmns}}$), and formation of a NS remnant (which could be a HMNS, SMNS or stable NS) that does not collapse within simulation time ($p_{\rm\textsc{nc}}$). Finally, Classifier C includes four classes: prompt collapse to BH ($p_{\rm\textsc{pcbh}}$), production of a HMNS that collapses during the simulation, with $2\,{\rm ms}<\tau_{\rm BH}<5\,{\rm ms}$ ($p_{\rm\textsc{short}}$) or $\tau_{\rm BH}>5\,{\rm ms}$ ($p_{\rm\textsc{long}}$), and formation of a remnant that does not collapse within the simulation time ($p_{\rm\textsc{nc}}$).